# Robust Inference for Non-Linear Regression Models with Applications in Enzyme Kinetics

### v1

@@ 3 Robust Wald-type tests based on the MDPDE
Basu et al. [7] first defined and studied a generalized class of robust Wald-type tests based on the MDPDE under the IID setup, which was later extended for the general INH setup in Basu et al. [5]. Here we will use the same approach to develop robust tests of hypotheses for our NLR models.
@@ 3.1 General Wald-type tests of NLR model parameters
where $\bm{L}$ is a given $r\times p$ matrix with rank$(\bm{L})=r$ and $\bm{l}_{0}$ is a given $r$-vector. Recall that the full parameter vector is $\bm{\theta}=\left(\bm{\beta}^{T},\sigma^{2}\right)^{T}$ with the parameter space $\bm{\Theta}\subseteq\mathbb{R}^{p}\times\mathbb{R}^{+}$. The restricted parameter space under the null hypothesis in (23) is given by
$\bm{\Theta_{0}}=\left\{\left(\bm{\beta}^{T},\sigma^{2}\right)^{T}:\bm{L\beta=l_{0}}\mbox{ and }0<\sigma^{2}<\infty\right\}.$
Also, the MDPDE of $\bm{\theta}=\left(\bm{\beta}^{T},\sigma^{2}\right)^{T}$ with tuning parameter $\alpha\geq 0$ has been denoted by $\bm{\widehat{\theta}_{\alpha}}=\left(\bm{\widehat{\beta}_{\alpha}}^{T},\widehat{\sigma}_{\alpha}^{2}\right)^{T}$. Then, we define the corresponding Wald-type test statistics for testing (23) as given by
$W_{n,\alpha}=\frac{1}{v_{1\alpha}\widehat{\sigma}_{\alpha}^{2}}\bm{\left(L\widehat{\beta}_{\alpha}-l_{0}\right)^{T}\left[L\left(\dot{\mu}\left(\widehat{\beta}_{\alpha}\right)^{T}\dot{\mu}\left(\widehat{\beta}_{\alpha}\right)\right)^{-1}L^{T}\right]^{-1}\left(L\widehat{\beta}_{\alpha}-l_{0}\right)}.$
Note that, under the null hypothesis in (23), one can easily show that, $W_{n,\alpha}$ asymptotically follows $\chi^{2}_{r}$, the chi-square distribution with $r$ degrees of freedom. So, the null hypothesis will be rejected at level of significance $\gamma$ if the observed value of $W_{n,\alpha}>\chi^{2}_{\gamma;r}$, the $(1-\gamma)$-th quantile of $\chi^{2}_{r}$ distribution.
@@
The above Wald-type tests coincide with the classical Wald-test at $\alpha=0$, due to the fact that the MDPDE coincides with the MLE at $\alpha=0$.
By the general theory discussed in Basu et al. [5], the above Wald-type test can be shown to be consistent at any fixed alternative for every values of $\alpha\geq 0$. Now, let us consider the contiguous alternative hypotheses of the form
@@
with $\bm{d}\in\mathbb{R}^{r}-\{\bm{0}_{r}\}$ be a fixed vector. Then, by Theorem 8 of Basu et al. [5], under $H_{1,n}$, $W_{n,\alpha}\xrightarrow{\mathcal{D}}\chi^{2}_{r}(\delta)$, the non-central chi-square distribution with degrees of freedom $r$ and the non-centrality parameter $\delta$, where
$\delta=\frac{1}{nv_{1\alpha}\sigma_{0}^{2}}\bm{d^{T}\left[L\left(\dot{\mu}(\beta_{0})^{T}\dot{\mu}(\beta_{0})\right)^{-1}L^{T}\right]^{-1}d},$ (25)
with $\left(\bm{\beta_{0}}^{T},\sigma_{0}^{2}\right)^{T}\in\bm{\Theta_{0}}$. Hence, an approximate expression of asymptotic contiguous power under $H_{1,n}$ in (24) is given by
$\pi_{n,\alpha}=1-G_{\chi^{2}_{r}(\delta)}\left(\chi^{2}_{\gamma;r}\right),$
where $G_{\chi^{2}_{r}(\delta)}(\cdot)$ denotes the distribution function of $\chi^{2}_{r}(\delta)$.
Let us now consider the setup of Section 2.3 in order to study the robustness of the proposed Wald-type test statistics. We first define the statistical functional corresponding to the test statistics $W_{n,\alpha}$ as given by
@@
Let $\bm{\theta_{0}}=\left(\bm{\beta_{0}}^{T},\sigma_{0}^{2}\right)^{T}\in\bm{\Theta}_{0}$ and $\bm{F_{\theta_{0}}}$ be the vector of the model distributions under the null hypotheses. It follows from the theory of general Wald-type tests (see, e.g., [7, 5]) that the first order influence function of the test functional $W_{\alpha}$ is zero when evaluated at $\bm{G}=\bm{F_{\theta_{0}}}$. The corresponding second order IF of the Wald-type test statistic $W_{\alpha}$ at $\bm{G}=\bm{F_{\theta_{0}}}$, when there is contamination only in the $i_{0}$-th observation, is obtained as
$IF_{i_{0}}^{(2)}\left(t_{i_{0}};W_{\alpha},\bm{F_{\theta_{0}}}\right)\\ =\frac{2}{nv_{1\alpha}\sigma_{0}^{2}}IF_{i_{0}}\left(t_{i_{0}};\bm{T^{\beta}}_{\alpha},\bm{F_{\theta_{0}}}\right)^{T}\bm{L^{T}\left[L\left(\dot{\mu}(\beta_{0})^{T}\dot{\mu}(\beta_{0})\right)^{-1}L^{T}\right]^{-1}L}IF_{i_{0}}\left(t_{i_{0}};\bm{T^{\beta}}_{\alpha},\bm{F_{\theta_{0}}}\right).$
Note that, this second order IF directly depends on the IF of the MDPDE functional $\bm{T^{\beta}}_{\alpha}$ of $\bm{\beta}$. Since $IF_{i_{0}}(t_{i_{0}};\bm{T^{\beta}}_{\alpha},\bm{F_{\theta_{0}}})$ is bounded for all $\alpha>0$ and unbounded at $\alpha=0$, so is $IF_{i_{0}}^{(2)}\left(t_{i_{0}};W_{\alpha},\bm{F_{\theta_{0}}}\right)$. This ensures the robustness of the Wald-type test for any $\alpha>0$ as well.
We can also study the robustness of the level and power of these Wald-type tests by examining its level influence function (LIF) and power influence function (PIF). From Theorem 12 of Basu et al. [5], we can conclude that the LIF of the MDPDE based Wald-type test is identically zero whenever the IF of the MDPDE used in constructing the test is bounded, i.e., for all $\alpha>0$. Further, its PIF at level of significance $\gamma$ is given by
$PIF\left(\bm{t};W_{\alpha},\bm{F_{\theta_{0}}}\right)=K_{r}^{*}(\delta)\frac{1}{nv_{1\alpha}\sigma_{0}^{2}}\bm{d^{T}\left[L\left(\dot{\mu}(\beta_{0})^{T}\dot{\mu}(\beta_{0})\right)^{-1}L^{T}\right]^{-1}}IF\left(\bm{t;T_{\alpha},F_{\theta_{0}}}\right),$
@@
3.2 Tests for scalar parameter against one sided alternative
Let us now consider the problem of testing a single parameter component, against a one-sided alternative, i.e., testing the hypothesis
@@ 3.2 Tests for scalar parameter against one sided alternative
where $\beta_{k}$ is the $k$-th component of regression parameter $\bm{\beta}$ in our NLR model (3). Under the both sided alternative $H_{1}^{*}:\beta_{k}\neq\beta_{k}^{0}$, the general MDPDE based Wald-type test statistics can be obtained from discussion of the previous subsection, which asymptotically follow $\chi^{2}_{1}$. However, for testing the one-sided alternative in (26), since we are dealing with a scalar parameter, we can use the test statistics
$\widetilde{W}_{n,\alpha}=\frac{\widehat{\beta}_{k\alpha}-\beta_{k}^{0}}{\sqrt{v_{1\alpha}\widehat{\sigma}_{\alpha}^{2}s_{k}}},$ (27)
@@
where $\widehat{\beta}_{k\alpha}$ is the $k$-th component of $\bm{\widehat{\beta}_{\alpha}}$, and $s_{k}$ is the $(k,k)$-th element of the matrix $\bm{\left(\dot{\mu}\left(\widehat{\beta}_{\alpha}\right)^{T}\dot{\mu}\left(\widehat{\beta}_{\alpha}\right)\right)}^{-1}$. It is evident from Theorem 2.1 that, under $H_{0}$ in (26), $\widetilde{W}_{n,\alpha}$ asymptotically follows a standard normal distribution. So, this null hypothesis will be rejected against the one sided alternative $H_{1}$ in (26), at level of significance $\gamma$, if the observed value of $W_{n,\alpha}$ exceeds $z_{\gamma}$, the $(1-\gamma)$-th quantile of $\mathcal{N}(0,1)$.
@@
Further, with respect to the one-sided contiguous alternative hypotheses of the form
$H_{1,n}:\beta_{k}=\beta_{k}^{0}+n^{-1/2}d,\ d>0,$
the large sample approximate expression of the power is given by
@@
$\widetilde{\pi}_{n,\alpha}=P_{H_{1,n}}\left[\widetilde{W}_{n,\alpha}>z_{\gamma}\right]\approx 1-\Phi\left(z_{\gamma}-\sqrt{\frac{1}{\sigma^{2}}\left(1+\frac{\alpha^{2}}{1+2\alpha}\right)^{-3/2}}\frac{d}{\sqrt{ns_{k}}}\right).$
@@
Note that, this asymptotic contiguous power $\widetilde{\pi}_{n,\alpha}$ depends on the covariates and the model mean function only through the quantity $d^{*}=\frac{d}{\sqrt{ns_{k}}}$. Table 2 shows the empirical values of $\widetilde{\pi}_{n,\alpha}$ at 5% level of significance, over different values of tuning parameter $\alpha$ and for different values of $d^{*}$, with $\sigma^{2}=1$. Clearly, for fixed $\alpha$, the power increases with $d^{*}$; in fact $d^{*}=0$ gives the level of the test. But, for fixed $d^{*}>0$, the power of our proposed test decreases with increasing values of $\alpha$, which is quite natural, following the pattern of the ARE of the MDPDE used to construct these tests. Additionally, the loss in power is not quite significant compared to the classical MLE based test (at $\alpha=0$) for small positive values of $\alpha$ (and also for larger $\alpha$ values when $d^{*}$ is large).
Table 2: Asymptotic contiguous power of the Wald-type test for testing $H_{0}$ in (26) against $H_{1,n}$
$d^{*}$ $\alpha$
0 0.1 0.3 0.5 0.7 1
0 0.050 0.050 0.050 0.050 0.050 0.050
1 0.260 0.258 0.247 0.233 0.219 0.201
2 0.639 0.634 0.608 0.574 0.538 0.487
3 0.912 0.909 0.891 0.865 0.833 0.780
5 1.000 1.000 0.999 0.998 0.997 0.991
7 1.000 1.000 1.000 1.000 1.000 1.000
@@ 4 Application to the Michaelis-Menten (MM) Model
where $\mu(x_{i},\bm{\beta})=\frac{\beta_{1}x_{i}}{\beta_{2}+x_{i}}$ with $\bm{\beta}=(\beta_{1},\beta_{2})^{T}=(V_{max},K_{m})^{T}$ being the regression parameter vector and the random errors $\epsilon_{i}\sim\mathcal{N}(0,\sigma^{2})$, independently for all $i=1,\ldots,n$. As the mean function here is not defined at $\beta_{2}=-x_{i}$, for any $i$, so we must take the parameter space for $\beta_{2}$ to be a subset of $\mathbb{R}\setminus\{-x_{1},\ldots,-x_{n}\}$; in applications, we generally have $\beta_{2}>0$, which ensures that the corresponding parameter space ($\mathbb{R}^{+}$) is independent of the values of the covariate, and hence the resulting mean function always remains well-defined. The parameter space for $\beta_{1}$ may be taken as $\mathbb{R}$.
For the MM model (28), let us denote $\widehat{\bm{\theta}}_{\alpha}=\left(\widehat{\bm{\beta}}_{\alpha}^{T},\widehat{\sigma}_{\alpha}^{2}\right)^{T}=\left(\widehat{\beta_{1}}_{\alpha},\widehat{\beta_{2}}_{\alpha},\widehat{\sigma}_{\alpha}^{2}\right)^{T}$ to be the MDPDE of $\bm{\theta}=\left(\bm{\beta}^{T},\sigma^{2}\right)^{T}$ corresponding to a tuning parameter $\alpha\geq 0$, defined as the minimizer of the objective function in (5) with the particular model specific form of $\mu(x_{i},\bm{\beta})=\frac{\beta_{1}x_{i}}{\beta_{2}+x_{i}}$. The corresponding estimating equations are given by (6) and (7), where we now have $\nabla_{\bm{\beta}}\mu(\bm{x_{i}},\bm{\beta})=\left(\frac{x_{i}}{\beta_{2}+x_{i}},\ -\frac{\beta_{1}x_{i}}{(\beta_{2}+x_{i})^{2}}\right)^{T}$.

### v2

@@ 3 Robust Wald-type tests based on the MDPDE
Basu et al. [7] first defined and studied a generalized class of robust Wald-type tests based on the MDPDE under the IID setup, which was later extended for the general INH setup in Basu et al. [5]. Here, we use the same approach to develop robust tests of hypotheses for our NLR models.
@@ 3.1 General Wald-type tests of NLR model parameters
[…]
$W_{n,\alpha}=\frac{1}{v_{1\alpha}\widehat{\sigma}_{\alpha}^{2}}\bm{\left(L\widehat{\beta}_{\alpha}-l_{0}\right)^{T}\left[L\left(\dot{\mu}\left(\widehat{\beta}_{\alpha}\right)^{T}\dot{\mu}\left(\widehat{\beta}_{\alpha}\right)\right)^{-1}L^{T}\right]^{-1}\left(L\widehat{\beta}_{\alpha}-l_{0}\right)},$
where $v_{1\alpha}$ is as defined in Theorem 2.1. Note that, under the null hypothesis in (22), one can easily show that, $W_{n,\alpha}$ asymptotically follows $\chi^{2}_{r}$, the chi-square distribution with $r$ degrees of freedom. So, the null hypothesis will be rejected at level of significance $\gamma$ if the observed value of $W_{n,\alpha}>\chi^{2}_{\gamma;r}$, the $(1-\gamma)$-th quantile of $\chi^{2}_{r}$ distribution.
@@
The above Wald-type test coincides with the classical Wald-test at $\alpha=0$, due to the fact that the MDPDE coincides with the MLE at $\alpha=0$.
By the general theory discussed in Basu et al. [5], the above Wald-type test can be shown to be consistent at any fixed alternative for every value of $\alpha\geq 0$. Now, let us consider the contiguous alternative hypotheses of the form
@@
with $\bm{d}\in\mathbb{R}^{r}\setminus\{\bm{0}_{r}\}$ be a fixed vector. Then, by Theorem 8 of Basu et al. [5], under $H_{1,n}$, $W_{n,\alpha}\xrightarrow{\mathcal{D}}\chi^{2}_{r}(\delta)$, the non-central chi-square distribution with degrees of freedom $r$ and the non-centrality parameter $\delta$, where
$\delta=\frac{1}{v_{1\alpha}\sigma_{0}^{2}}\bm{d}^{T}\left[\bm{L}\bm{M}^{-1}(\bm{\beta}_{0})\bm{L}^{T}\right]^{-1}\bm{d},\mbox{ with }\bm{M}(\bm{\beta}_{0})=\lim_{n\to\infty}\frac{1}{n}\bm{\dot{\mu}}(\bm{\beta}_{0})^{T}\bm{\dot{\mu}}(\bm{\beta}_{0}),~~\left(\bm{\beta}_{0}^{T},\sigma_{0}^{2}\right)^{T}\in\bm{\Theta}_{0}.$~~ (24)
Hence, an approximate expression of asymptotic contiguous power under $H_{1,n}$ in (23) is given by $\Pi_{\alpha}=1-G_{\chi^{2}_{r}(\delta)}\left(\chi^{2}_{\gamma;r}\right)$, where $G_{\chi^{2}_{r}(\delta)}(\cdot)$ denotes the distribution function of $\chi^{2}_{r}(\delta)$.
3.2 Influence functions of the Wald-type tests
Let us now consider the setup of Section 2.3 to study the robustness of the proposed Wald-type test statistics. We first define the statistical functional corresponding to the test statistics $W_{n,\alpha}$ as
@@ 3.2 Influence functions of the Wald-type tests
Let $\bm{\theta}_{0}=\left(\bm{\beta}_{0}^{T},\sigma_{0}^{2}\right)^{T}\in\bm{\Theta}_{0}$ and $\bm{F}_{\bm{\theta}_{0}}$ be the vector of the model distributions under the null hypotheses. It follows from the theory of general Wald-type tests (see, e.g., [7, 5]) that the first-order IF of the test functional $W_{\alpha}$ (as defined for the estimators earlier) is zero when evaluated at $\bm{G}=\bm{F}_{\bm{\theta}_{0}}$, and so we need to consider the second-order IF.
The second-order IF of any statistical functional $W$, at the vector of distribution functions $\bm{G}$, with contamination only at the $i$-th observation at the contamination point $t_{i}$ is defined as
$IF_{i}^{(2)}(t_{i};W,\bm{G})=\frac{\partial^{2}}{\partial\epsilon^{2}}W(G_{1},\ldots,G_{i-1},G_{i,\epsilon},G_{i+1},\ldots,G_{n})\bigg|_{\epsilon=0},$
where $G_{i,\epsilon}=(1-\epsilon)G_{i}+\epsilon\Delta_{t_{i}}$, with $\Delta_{t_{i}}$ being the one-point distribution supported on $\{t_{i}\}$ (see, e.g., [16, 5]). Intuitively, it measures the second-order bias approximation caused by infinitesimal contamination at the outlying point. For the present Wald-type test functional $W_{\alpha}$, this second-order IF at $\bm{G}=\bm{F}_{\bm{\theta}_{0}}$ turns out to be
$IF_{i_{0}}^{(2)}\left(t_{i_{0}};W_{\alpha},\bm{F}_{\bm{\theta}_{0}}\right)=\frac{2}{nv_{1\alpha}\sigma_{0}^{2}}IF_{i_{0}}\left(t_{i_{0}};\bm{T^{\beta}}_{\alpha},\bm{F}_{\bm{\theta}_{0}}\right)^{T}\bm{L}^{T}\left[\bm{L}\left(\bm{\dot{\mu}}(\bm{\beta}_{0})^{T}\bm{\dot{\mu}}(\bm{\beta}_{0})\right)^{-1}\bm{L}^{T}\right]^{-1}\bm{L}~IF_{i_{0}}\left(t_{i_{0}};\bm{T^{\beta}}_{\alpha},\bm{F}_{\bm{\theta}_{0}}\right),$
when there is contamination only in the $i_{0}$-th observation. Note that, this second-order IF directly depends on the IF of the MDPDE functional $\bm{T^{\beta}}_{\alpha}$ of $\bm{\beta}$. Since $IF_{i_{0}}(t_{i_{0}};\bm{T^{\beta}}_{\alpha},\bm{F}_{\bm{\theta}_{0}})$ is bounded for all $\alpha>0$ and unbounded at $\alpha=0$, so is $IF_{i_{0}}^{(2)}\left(t_{i_{0}};W_{\alpha},\bm{F}_{\bm{\theta}_{0}}\right)$. This ensures the robustness of the Wald-type test for any $\alpha>0$ as well.
We can also study the robustness of the level and power of these Wald-type tests by examining their level influence function (LIF) and power influence function (PIF). A brief discussion of the LIF and PIF under general INH set-ups (including their definitions) and for the Wald-type test statistics is provided in Web Appendix A.3 for the sake of completeness. From Equations (9)-(10) presented there (or, directly from Theorem 12 of Basu et al. [5]), we can conclude that the LIF of the MDPDE-based Wald-type test is identically zero whenever the IF of the MDPDE used in constructing the test is bounded, i.e., for all $\alpha>0$. Further, its PIF at level of significance $\gamma$ is given by
$PIF\left(\bm{t};W_{\alpha},\bm{F}_{\bm{\theta}_{0}}\right)=K_{r}^{*}(\delta)\frac{1}{nv_{1\alpha}\sigma_{0}^{2}}\bm{d}^{T}\left[\bm{L}\left(\bm{\dot{\mu}}(\bm{\beta}_{0})^{T}\bm{\dot{\mu}}(\bm{\beta}_{0})\right)^{-1}\bm{L}^{T}\right]^{-1}IF\left(\bm{t};\bm{T}_{\alpha},\bm{F}_{\bm{\theta}_{0}}\right),$
@@
3.3 Tests for scalar parameter against one-sided alternative
Let us now consider the problem of testing a single parameter component against a one-sided alternative, i.e., testing the hypothesis
@@ 3.3 Tests for scalar parameter against one-sided alternative
where $\beta_{k}$ is the $k$-th component of regression parameter $\bm{\beta}$ in our NLR model (3). Under the both-sided alternative $H_{1}^{*}:\beta_{k}\neq\beta_{k}^{0}$, the general MDPDE-based Wald-type test statistics can be obtained from the discussion of the previous subsection, which asymptotically follow $\chi^{2}_{1}$. However, for testing the one-sided alternative in (25), since we are dealing with a scalar parameter, we can use the test statistic
$\widetilde{W}_{n,\alpha}=\frac{\widehat{\beta}_{k\alpha}-\beta_{k}^{0}}{\sqrt{v_{1\alpha}\widehat{\sigma}_{\alpha}^{2}s_{kk}}},$ (26)
@@
where $\widehat{\beta}_{k\alpha}$ is the $k$-th component of $\bm{\widehat{\beta}}_{\alpha}$, and $s_{kk}$ is the $(k,k)$-th element of the matrix $(\bm{\dot{\mu}}(\widehat{\bm{\beta}}_{\alpha})^{T}\bm{\dot{\mu}}(\widehat{\bm{\beta}}_{\alpha}))^{-1}$. It is evident from Theorem 2.1 that, under $H_{0}$ in (25), $\widetilde{W}_{n,\alpha}$ asymptotically follows a standard normal distribution. So, this null hypothesis will be rejected against $H_{1}$ in (25), at level of significance $\gamma$, if the observed value of $W_{n,\alpha}$ exceeds $z_{\gamma}$, the $(1-\gamma)$-th quantile of $\mathcal{N}(0,1)$.
@@
Further, one can observe that, under the one-sided contiguous alternative hypotheses of the form $H_{1,n}:\beta_{k}=\beta_{k}^{0}+n^{-1/2}d,\ d>0$, we have
@@
$\frac{\left(\widehat{\beta}_{k\alpha}-\beta_{k}^{0}-n^{-1/2}d\right)}{\sqrt{v_{1\alpha}\widehat{\sigma}_{\alpha}^{2}s_{kk}}}\xrightarrow[]{\mathcal{D}}\mathcal{N}(0,1),\mbox{ and hence }~\left[\widetilde{W}_{n,\alpha}-\frac{d}{\sqrt{v_{1\alpha}\widehat{\sigma}_{\alpha}^{2}ns_{kk}}}\right]\xrightarrow[]{\mathcal{D}}\mathcal{N}(0,1).$
Thus, the asymptotic contiguous power of the test based on $\widetilde{W}_{n,\alpha}$ is given by
$\widetilde{\Pi}_{\alpha}=\lim_{n\to\infty}P_{H_{1,n}}\left[\widetilde{W}_{n,\alpha}>z_{\gamma}\right]=1-\Phi\left(z_{\gamma}-\left(1+\frac{\alpha^{2}}{1+2\alpha}\right)^{-3/4}\frac{d^{*}}{\sigma}\right),$ (27)
@@
where $d^{*}=\frac{d}{\sqrt{s_{kk}^{*}}}$ with $s_{kk}^{*}=\lim_{n\to\infty}ns_{kk}$. Note that this asymptotic contiguous power $\widetilde{\Pi}_{\alpha}$ depends on the covariates and the model mean function only through the quantity $d^{*}$. Table C.2 of Web Appendix C shows the empirical values of $\widetilde{\Pi}_{\alpha}$ at 5% level of significance, over different values of tuning parameter $\alpha$ and for various values of $d^{*}$ (which may be obtained based on suitable covariate values and $d$), with $\sigma=1$. Clearly, for fixed $\alpha$, the power increases with $d^{*}$; in fact, $d^{*}=0$ gives the level of the test. But, for fixed $d^{*}>0$, the power of our proposed test decreases with increasing values of $\alpha$, which is quite natural, following the pattern of the ARE of the MDPDE used to construct these tests. Additionally, the loss in power is not quite significant compared to the classical MLE-based test (at $\alpha=0$) for small positive values of $\alpha$ (and also for larger $\alpha$ values when $d^{*}$ is large).
@@ 4 Application to the Michaelis-Menten (MM) Model
where $\mu(x_{i},\bm{\beta})=\frac{\beta_{1}x_{i}}{\beta_{2}+x_{i}}$ with $\bm{\beta}=(\beta_{1},\beta_{2})^{T}=(V_{max},K_{m})^{T}$ being the regression parameter vector and the random errors $\epsilon_{i}\sim\mathcal{N}(0,\sigma^{2})$, independently for all $i=1,\ldots,n$. As the mean function here is not defined at $\beta_{2}=-x_{i}$, for any $i$, so we must take the parameter space for $\beta_{2}$ to be a subset of $\mathbb{R}\setminus\{-x_{1},\ldots,-x_{n}\}$; in applications, we generally have $\beta_{2}>0$, which ensures that the corresponding parameter space ($\mathbb{R}^{+}$) is independent of the values of the covariate, and hence the resulting mean function always remains well-defined. The parameter space for $\beta_{1}$ may be taken as $\mathbb{R}$. For the MM model (28), let us denote $\widehat{\bm{\theta}}_{\alpha}=(\widehat{\bm{\beta}}_{\alpha}^{T},\widehat{\sigma}_{\alpha}^{2})^{T}=(\widehat{\beta}_{1\alpha},\widehat{\beta}_{2\alpha},\widehat{\sigma}_{\alpha}^{2})^{T}$ to be the MDPDE of $\bm{\theta}=\left(\bm{\beta}^{T},\sigma^{2}\right)^{T}$ corresponding to a tuning parameter $\alpha\geq 0$, defined as the minimizer of the objective function in (5) with the particular model specific form of $\mu(x_{i},\bm{\beta})=\frac{\beta_{1}x_{i}}{\beta_{2}+x_{i}}$. The corresponding estimating equations are given by (6) and (7), where we now have $\nabla_{\bm{\beta}}\mu(\bm{x}_{i},\bm{\beta})=\left(\frac{x_{i}}{\beta_{2}+x_{i}},\ -\frac{\beta_{1}x_{i}}{(\beta_{2}+x_{i})^{2}}\right)^{T}$.